OpenStudy (anonymous):

What is the standard form of (5 + 2i)(5 – 2i)?

OpenStudy (trojanpoem):

25 - 4i^2 25 + 4 = 29

OpenStudy (anonymous):

What is the product of (1 + i)(2 + i)?

OpenStudy (trojanpoem):

multiply each element 2+ i + 2i + i^2 2 + 3i - 1 1 + 3i

OpenStudy (anonymous):

Simplify the expression 5^3 × 5^8.

OpenStudy (trojanpoem):

in multiplication you sum the powers 5^(3+8)= 5^11

OpenStudy (anonymous):

Find the difference: (6y^3 + 17y − 3) − (4y^3 − 11y + 9)

OpenStudy (trojanpoem):

Direct summation 2y^3 + 28 y - 12

OpenStudy (trojanpoem):

I don't think , that helps, I am answering not you.

OpenStudy (anonymous):

Find the sum: (3x^2 + 5x − 8) + (5x^2 − 13x − 5)

OpenStudy (trojanpoem):

8 x^2 - 8 x - 13

OpenStudy (anonymous):

Simplify the expression (2x − 9)(x + 6).

OpenStudy (trojanpoem):

2x^2 +6x - 9 x - 54 2x^2 - 3x - 54

OpenStudy (anonymous):

Simplify the expression 4x(2x^2 − 7x + 3).

OpenStudy (anonymous):

?

OpenStudy (trojanpoem):

8x^3 - 28 x^2 + 12x